OpenStudy (anonymous):

Air Is being pumped into a sphere balloon so that its volume increase a rate of 20 cm^3/s. How fast is radius of the balloon increasing when the diameter is 10 cm? V= 3/4 pi*r^2*h

OpenStudy (anonymous):

This is a related rates problem... you first need to decide what the "snapshot" variables are. That means, what are the starting measurements. THEN, you need to determine the "dynamic" variables. This means, what quantities are changing over time. Then, you need to differentiate the Volume equation (keeping in mind that you must differentiate EACH variable with respect to time). Try that and see if it helps you... if not, let me know.

OpenStudy (anonymous):

so d/dt= 3/4 2r (r) h v=20 cm^3 diameter is 10=5 like this ??

OpenStudy (anonymous):

dV/dt=20

OpenStudy (anonymous):

dr/dt=?

OpenStudy (anonymous):

forget to put V'

OpenStudy (anonymous):

dr/dt=5?

OpenStudy (anonymous):

dr/dt = the rate at which your radius is changing

OpenStudy (anonymous):

which is what the question is asking

OpenStudy (anonymous):

then i need find the derivate of r?

OpenStudy (anonymous):

no you need to first find a proportion that relates height and radius

OpenStudy (anonymous):

found out that i wrote the wrong formula for the volume of the sphere

OpenStudy (anonymous):

4/3 r^3 pi

OpenStudy (anonymous):

correct alright so now that you have that you need to use implicit differentiation

OpenStudy (anonymous):

differentiate in respects to t. what do you get?

OpenStudy (anonymous):

\[\frac{dV}{dt}=\frac{4\pi\frac{d}{dt}[r^3]}{3}\]

OpenStudy (anonymous):

2-= 4pi (5)^2*r' or dr/dt

OpenStudy (anonymous):

don't plug in anything before you differentiate because the rat is always changign

OpenStudy (anonymous):

k

OpenStudy (anonymous):

The volume expression for a spherical balloon is:\[V=\frac{4 \pi r^3}{3} \]The Total Derivative of the above is:\[dV=4 \pi r^2 dr \]Plug in the given values for dV and dr and then solve for dr.

OpenStudy (anonymous):

that wat i did @robtobey

OpenStudy (anonymous):

\[20=4pi(5)^2(\frac{dr}{dt})\]= \[\frac{20}{4\pi(5)^2}=\frac{dr}{dt}\]

OpenStudy (anonymous):

dr/dt=1/5 pi

OpenStudy (anonymous):

thank for the help